Paul is an 11-month-old who you want to prescribe a corticosteroid. His weight is 14 pounds, 6 ounces. You will prescribe Pediapred 6.7 mg/5ml. You wish to give 2 mg/kg/day in three divided dosages for 5 days. Please calculate the dosage, show your work.
Answer #1

Given

Order:Pediapred

Dose:6.7mg

Available:6.7mg/5ml

Disease to be given 2mg/kg/day

Weight 14 pounds, 6 ounces

To convert to kg multiply by 0.454

=14.6×0.454

=6.6kg

As 1kg=2mg/day

Then 6.6 kg=??mg/day

Cross multiply and divide

=6.6×2÷1

=6.6×2

=13.2mg/day

Frequency:Three doses in a day

Therefore to calculate one dose 13.2÷3= 4.4 mg

So 1 dose=4.4 ml

As availability is 6.7mg/5ml

Therefore 4.4 mg=??ml

Cross multiply and divide

=4.4×5÷6.7

=4.4×0.7

=3.08 or 3ml

Ans:A dosage of 4.4 mg will be given per dose by administering 3ml